The Ksp of BaCO3 is 5.0 x 10-9. What is the soliubility of barium carbonate in a solution that contains 1.18 M of Na2CO3?
Answer #1

Consider dissolution of BaCO3 in water as BaCO3(s) media%2Fabe%2Fabe16a14-4c4c-4df9-bb16-e4 Ba 2+(aq) + CO32-(aq)

K sp = [Ba 2+ ] [CO32-] = 5.0 x 10 -09  

Consider S be the solubility of BaCO3 in mol / L , then [Ba 2+ ] =[CO32-]= S mol / L.

But solution contain 1.18 M Na2CO3 hence [CO32-] = 1.18 + S.

K sp = [Ba 2+ ] [CO32-] = 5.0 x 10 -09  

S x ( 1.18+S) = 5.0 x 10 -09  

Here S is very small as compared to 1.18 hence we can write 1.18 + S = 1.18

S x ( 1.18) = 5.0 x 10 -09  

S = 5.0 x 10 -09 / 1.18

S = 4.237 x 10 -09 mol / L

ANSWER: Solubility of BaCO3 in water = 4.237 x 10 -09 mol / L